What is the return on equity for 2017?
the following information to answer this question. Bayside, Ind. 2017 Income Statement ( in thousands) $6,190 4,650 445 et sales Cost of goods sold Depreciation Earnings before interest and taxes Interest paid Taxable income Taxes Net income $1,095 40 $1,055 317 $ 738 Bayside, Inc. 2016 and 2017 Balance Sheets ($ in thousands) 2016 2017 $1,805$1,825 900700 3,420 3,370 970 1,220 2016 2017 s 160 265 Accounts payable Cash Accounts rec Inventory Total Net fixed assets Total assets 1,170 1,010 Long-term debt 1,885 2,130 Conmon stock $3,215 $3,405 Retained earnings 3.880 3,710 $7,095 $7,115 Total liab. & equity $7,095$7,115 What is the return on equity for 2017?
Answer #1

Return on equity = (Net income / Average equity) × 100

Given, Net income = $738 thousand

Average equity = [Beginning + Ending] / 2

                        = [(3,420 + 970) + (3,370 + 1,220)] / 2

                        = (4,390 + 4,590) / 2

                        = 8,980 / 2

                        = $4,490 thousand

Return on equity = (Net income / Average equity) × 100

                            = (738 / 4490) × 100

                            = 16.436 ….

                            = 16.44% (Answer)